# Synth-to-Real Unsupervised Domain Adaptation for Instance Segmentation

### v2

@@ 4 Experiments
4.1 Datasets
In the experiments, we set Cityscapes [8] as the target domain, and UrbanSyn [15], Synscapes [66] and SYNTHIA [50] as source domains. We conduct experiments in three synth-to-real domain settings: UrbanSyn Cityscapes, Synscapes Cityscapes, and SYNTHIA Cityscapes.
Cityscapes is a street-scene dataset that contains 5000 images (2,975, 500, and 1,525 images for training, validation, and testing, respectively) and their corresponding labels for 19 semantic classes and 8 instance classes. Cityscapes is a commonly used real-domain dataset with high-resolution () images in the field of autonomous driving.
UrbanSyn is a high-quality synthetic dataset tailored to urban driving scenarios. It comprises a total of 7,539 images with a resolution of pixels. The images in this dataset have highly realistic rendering effects, making them closely resemble real-world scenes. UrbanSyn has the same instance and semantic classes as Cityscapes.
Synscapes is another physic-based-rendered urban scenes dataset, which includes 25,000 images at a resolution of . The labeled data covers 8 instance classes, but the division is different from Cityscapes. To make them consistent with Cityscapes, we reclassify the ‘Motorcyclist’ and ‘Bicyclist’ instances into ‘Motorcycle’, ‘Bicycle’, and ‘Rider’ leveraging semantic labels. Additionally, we filter the instances with high occlusion, high truncated, or a tiny area during training.
SYNTHIA consists of photo-realistic images in a resolution of rendered from a virtual city, with a subset compatible with Cityscapes. This subset contains 9,400 images with pixel-level semantic annotations. SYNTHIA covers 6 classes of instance annotation in Cityscapes, missing ‘Train’ and ‘Truck’.
@@ 4.2.1 Training
We adopt Mask2Former [6] with Swin Transformer [38] backbone pre-trained on COCO [36] as our base network . To minimize the color discrepancy between the source and target domains, we adopt the same image pre-processing as [20], aligning the style of source domain images to the target one by shifting the distribution of pixel values in CIELAB color space. For image processing, we apply the same data augmentation (flipping, random cropping, etc.) as [6] and set the image crop size to . The maximum number of samples N for the rare-class pool is set to 10, and the pseudo-label confidence threshold is 0.9. We set the numbers of iterations for source domain training to 10k, while the mixed data training is 20k iterations. For each stage of training, we set a batch size =3 considering optimal memory usage. We use AdamW [41] optimizer with an initial learning rate of 0.0001 and a weight decay of 0.05, as well as the learning rate warm-up and the ‘poly’ learning rate schedule. The hyper-parameters of loss , , and are set to 5.0, 5.0, 2.0 respectively. For three synth-to-real settings, we set the weights of bidirectional data mixing loss , to: 0.4, 0.6 for UrbanSyn Cityscapes and SYNTHIA Cityscapes, 0.75, 0.25 for Synscpase Cityscapes. Our model is trained with 3 NVIDIA A40 GPUs, each with 40GB of memory.
@@ 4.2.2 Evaluation
We evaluate the UDA performance of 8-class instance segmentation on the validation set of Cityscapes. Following the conventional evaluation criteria of instance segmentation, we take AP (Average Precision) and AP50 (Average Precision at IOU=0.5) for each class as our metrics for comparison, and compute the mAP (mean Average Precision) and mAP50 (mean Average Precision at IOU=0.5) for all classes. All the numbers present in the tables are in %.
@@ 4.3 Experimental Results
In Table 1, we summarize the results of models tested on the Cityscapes validation set. The table is categorized into two sections showing AP (left) and mAP (right) results, with each section organized from top to bottom into three synth-to-real settings. In all settings, we show results of the UDA4Inst model and the original instance segmentation baseline model Mask2Former [6] trained only on the source dataset. The purpose is to compare the performance difference in domain adaptation when training the same network with or without our proposed UDA method.
| AP | AP50 | |||||||||||||||||
| Model | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP50 |
| upper bound | 41.1 | 32.4 | 61.4 | 44.8 | 67.5 | 50.7 | 26.8 | 26.6 | 43.9 | 75.4 | 71.3 | 87.2 | 58.7 | 80.3 | 76.1 | 58.1 | 65.1 | 71.5 |
| UrbanSyn Cityscapes | ||||||||||||||||||
| w/o UDA | 25.1 | 19.5 | 39.8 | 37.6 | 51.5 | 46.7 | 19.2 | 16.2 | 32.0 | 52.0 | 48.4 | 61.5 | 48.6 | 66.2 | 70.0 | 42.6 | 46.0 | 54.4 |
| UDA4Inst | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 | 60.1 | 60.7 | 75.1 | 56.4 | 78.0 | 77.8 | 48.7 | 53.6 | 63.8 |
| Synscapes Cityscapes | ||||||||||||||||||
| w/o UDA | 27.2 | 19.2 | 39.5 | 24.0 | 46.7 | 33.1 | 18.7 | 16.6 | 28.1 | 53.9 | 44.3 | 67.8 | 31.9 | 60.6 | 58.1 | 39.9 | 45.4 | 50.2 |
| UDA4Inst | 30.3 | 27.1 | 47.5 | 34.5 | 58.5 | 46.3 | 22.3 | 19.1 | 35.7 | 59.9 | 63.2 | 77.9 | 45.3 | 72.8 | 63.5 | 46.1 | 51.7 | 60.1 |
| SYNTHIA Cityscapes | ||||||||||||||||||
| Zhang [70] | 9.4 | 3.8 | 22.2 | - | 23.5 | - | 3.2 | 2.4 | 10.8 | 28.8 | 17.8 | 39.9 | - | 37.5 | - | 12.8 | 11.9 | 24.8 |
| CVRN [27] | - | - | - | - | - | - | - | - | - | 34.4 | 25.3 | 38.7 | - | 38.1 | - | 10.1 | 8.7 | 25.9 |
| EDAPS [52] | 23.5 | 13.4 | 27.5 | - | 23.6 | - | 5.8 | 0.1 | 15.7 | 53.9 | 40.6 | 50.5 | - | 38.8 | - | 21.1 | 0.7 | 34.3 |
| w/o UDA | 26.3 | 15.2 | 38.3 | - | 43.0 | - | 13.6 | 10.9 | 24.6 | 53.4 | 40.8 | 68.1 | - | 56.5 | - | 36.4 | 36.1 | 48.6 |
| UDA4Inst | 28.1 | 19.6 | 47.9 | - | 58.7 | - | 18.5 | 15.0 | 31.3 | 57.7 | 52.5 | 76.9 | - | 68.6 | - | 42.6 | 43.3 | 56.9 |
(a) AP
(b) AP50
@@
To establish an upper bound, we retrain the Mask2Former model on the Cityscapes training set, reproducing similar validation results reported in their paper. It reflects the instance segmentation performance of models trained and tested on the target domain, i.e. without crossing domains. Overall, the first observation is that in all synth-to-real settings, UDA4Inst exhibits a significant improvement in both mAP and mAP50 compared to the baseline model. The most evident improvement is reflected in the setting of Synscapes Cityscapes, where the mAP and mAP50 increased by +7.6 and +9.9 points respectively. Our model achieves the highest 39.0 mAP in UrbanSyn Cityscapes, 35.7 mAP in Synscapes Cityscapes, and 31.3 mAP in SYNTHIA Cityscapes, which are very competitive. Among three settings, we observe that both UDA4Inst and the baseline show better adaptation performance in UrbanSyn Cityscapes. We hypothesize that these results reflect a smaller domain gap between UrbanSyn and Cityscapes.
In particular, UDA4Inst demonstrates consistent enhancement over the baseline model across all the classes in all synth-to-real settings, showcasing the strong efficacy of our proposed UDA method. To intuitively demonstrate the performance difference between these two models, we provide bar charts of class-wise AP and AP50 results in Fig. 7. We can see the gains of UDA4Inst over the baseline one. Notably, in UrbanSyn Cityscapes setting, UDA4Inst achieves 58.1 AP and 77.8 AP50 on the class of ‘Train’, which surpasses the upper bound model with gains of +7.4 and +1.7. It is also worth noting that for the classes of ‘Truck’ and ‘Bus’, UDA4Inst shows very close AP and AP50 results to the upper bound.
As a reference, at the bottom of the table, we also report results from Zhang [70], CVRN [27], and EDAPS [52] that were previously proposed in recent years to address UDA for instance segmentation in SYNTHIA Cityscapes setting. Overall, our proposed UDA4Inst outperforms the other relevant work with a significant improvement, specifically shown as gains of +20.5, +15.6 mAP compared to Zhang and EDAPS, as well as gains of +32.1, +31, +22.6 mAP50 compared to Zhang, CVRN, and EDAPS respectively. Note that we are the first to conduct experiments in the synth-to-real settings of UrbanSyn Cityscapes and Synscapes Cityscapes for instance segmentation.
To further qualitatively demonstrate the performance, in Fig. 8, we visualize the instance segmentation predictions and the error maps of the EDAPS, the baseline model, and UDA4Inst in the SYNTHIA Cityscapes setting. UDA4Inst shows the best segmentation performance, reflected in the fewest error predictions on objects such as cars, riders, and persons. This is consistent with the conclusion drawn from the results in Table 1.
4.3.2 Bidirectional Data Mixing.
The results in the UrbanSyn Cityscapes setting are shown in Table 2. The first row presents the results of a baseline model without applying data mixing. The mAP results reveal both S2T and T2S data mixing contribute to a better segmentation performance. T2S is slightly inferior to S2T, which we believe is because the imperfect generated target pseudo-labels can inevitably cause interference in data mixing. Not surprisingly, when combining S2T and T2S, our approach achieves the best result of 37.18 mAP, +5.23 higher than the plain model without data mixing. Only applying S2T mixing neglects the instances in target images. Through bidirectional data mixing, we strike a balance between learning target features and mitigating the negative impact of incorrect target instances on the model. Then, we perform hyper-parameters searching for the weights of the BDM loss when applying all three UDA methods. The results are shown in Table 3. The best result yields at the setting of = 0.4 and = 0.6.
@@
To illustrate the advantages of our proposed Bidirectional Data Mixing (BDM) method compared to the unidirectional one, we conduct experiments on different data mixing strategies without applying Rare-class Balancing (RcB) and Category Module Training (CMT).
@@
As we mentioned, we perform CMT in the hope that introducing this inductive bias can guide the model to better learn the features in classes. To validate our conjecture, in the UrbanSyn Cityscapes setting, we perform CMT on two category groups, which are random compositions of classes (setting 2-5 in Table 4). We also conduct an experiment of full-class training for comparison (setting 1). Overall, we observe that all four CMT models outperform the full-class training one. The best set of category groups is setting 5, which achieves the highest improvement of +3.5 mAP and +5.0 mAP50 compared to the model without applying CMT. We take this group setting as our default setting to conduct the rest experiments.
@@
To inspect the impact of components of UDA4Inst, we provide an ablation study in the UrbanSyn Cityscapes setting. Specifically, our interest focuses on the three components: BDM, RcB, and CMT. In Table 5, we present mAP results for models that trained without, with some, or with all three UDA methods. The ablative models are named from M1 to M5 (from top to bottom) for illustration.
Overall, we find that the M1 trained without applying any UDA method obtains the lowest mAP of 32.0 among all models. On the contrary, the M6, which is trained with all three UDA methods, achieves the highest score of 39.0. When comparing M1 and M2, we see that applying CMT can effectively improve the model performance by +2.8 mAP gains. The conclusion is further strengthened through the comparison of M3 and M4, where we observe a mAP gain of +1.6 when applying CMT. The effect of BDM is reflected in the performance difference between M1 and M3, where we see that it helps to improve the segmentation results from 32.0 to 36.1 mAP (+4.1 gains). Moreover, under the same circumstance of using CMT, BDM helps to improve the segmentation precision in M4 within most of the classes compared to M2, which contributes to a better mAP (+2.9 gains). The RcB method, which is based on BDM and CMT, is applied in M5. We observe that M4, which applies BDM without RcB, results in a -1.3 lower mAP than M5. Focusing on the defined rare classes, we see that RcB helps M5 to obtain +10.8 and +1.8 AP gains to the M4 in ‘Train’ and ‘Motorcycle’, respectively, and yields the best APs among all the models. By comparing the gains of mAP in ablative models, we conclude that among the three methods, the BDM is mostly beneficial for domain adaptation.
@@ 5 Conclusion
In this paper, we present a model named UDA4Inst that aims to address the challenge of synth-to-real UDA for instance segmentation in the field of autonomous driving. We introduce the three novel and effective methods in our model which are instance-level bidirectional data mixing, rare-class balancing, and category module training. In a comprehensive evaluation of three synth-to-real domain settings UrbanSyn Cityscapes, Synscapes Cityscapes, and SYNTHIA Cityscapes, we show that UDA4Inst achieves significant improvements in instance segmentation compared to the baseline model Mask2Former. This is the first work to provide competitive AP results on Synscapes and the recently published synthetic dataset UrbanSyn. Our proposed UDA4Inst model yields also state-of-the-art results on SYNTHIA. We look forward to the proposed method will contribute to the development of UDA and can be widely served for future research.
5.0.1 Acknowledgments

### v3

@@ 4.1 Implementation Details
we conduct experiments in this work in three synth-to-real domain settings: UrbanSyn Cityscapes, Synscapes Cityscapes, and SYNTHIA Cityscapes.
@@ 4.1.1 Training
We adopt Mask2Former [7] with Swin Transformer backbone [32] as our base network . The model is pre-trained on COCO [30] dataset. To minimize the color discrepancy, we adopt the same image pre-processing as [17], aligning the style of source domain images to the target ones by shifting the distribution of pixel values in CIELAB color space. We apply the same data augmentation (flipping, random cropping, etc.) as [7] and set the image crop size to . The maximum number of samples N for the rare-class pool is set to 10, and the pseudo-label confidence threshold is 0.9. We set the numbers of iterations for source domain training to 15k, while the mixed data training is 25k iterations. During validation, we follow the same checkpoint selection protocol used in [21]. For each stage of training, we set a batch size =3 considering optimal memory usage. We use AdamW [34] optimizer with an initial learning rate of 0.0001 and a weight decay of 0.05, as well as the ‘poly’ learning rate schedule with the power of 1. The hyper-parameters of loss , , and are set to 5.0, 5.0, 2.0 respectively. For three synth-to-real settings, we set the weights of bidirectional data mixing loss , to: 0.4, 0.6 for UrbanSyn Cityscapes and SYNTHIA Cityscapes, 0.75, 0.25 for Synscapes Cityscapes. For these experiments, UDA4Inst runs on 3 NVIDIA A40 GPUs, each with 48GB of memory.
@@ 4.1.2 Evaluation
We evaluate the UDA performance of 8-class instance segmentation on the validation set of Cityscapes. Following the conventional evaluation criteria of instance segmentation, we take AP (Average Precision) and AP50 (Average Precision at IOU>0.5) and their mean values as our metrics for comparison. All the numbers present in the tables are in %.
(a) AP
(b) AP50
@@ 4.2 Experimental Results
In Table 1, we summarize the results of models tested on the Cityscapes validation set. In three synth-to-real settings, we show results of the UDA4Inst model and the source-only vanilla Mask2Former model [7]. To establish an upper bound, we retrain the Mask2Former model on the Cityscapes training set, i.e., using human-provided labels, reproducing similar results in their paper [7]. Overall, the first observation is that in all settings, UDA4Inst exhibits a significant improvement in both mAP and mAP50 compared to the source-only model. The most evident improvement is reflected in the setting of Synscapes Cityscapes, where the mAP and mAP50 increased by +7.6 and +9.9 points respectively. Our model achieves the highest mAP (39.0) in UrbanSyn Cityscapes, 35.7 mAP in Synscapes Cityscapes, and 31.3 mAP in SYNTHIA Cityscapes. Note that we are the first to conduct experiments in the first two settings for instance segmentation. We observe that models show better adaptation performance in UrbanSyn Cityscapes. Notably, in this setting, UDA4Inst achieves 58.1 AP and 77.8 AP50 on the class of ‘Train’, which surpasses the upper bound model with gains of +7.4 and +1.7. UDA4Inst shows very close AP and AP50 results to the upper bound for ‘Truck’ and ‘Bus’ classes. We hypothesize that these results reflect a smaller domain gap between UrbanSyn and Cityscapes. In Fig. 4, we visualize the predictions and the error maps of the source-only model and UDA4Inst for this setting. We provide more visualizations in the supplementary material. In particular, UDA4Inst demonstrates consistent enhancement over the source-only model across all the classes in all synth-to-real settings. We provide bar charts of class-wise AP and AP50 results in Fig. 3 to see the gains of UDA4Inst over the source-only one.
@@
As a reference, at the bottom of the table, we list results from Zhang [55], CVRN [22], and EDAPS [44] that were proposed in recent years to address UDA for instance segmentation in SYNTHIA Cityscapes setting. Considering that the listed models used different base networks (i.e. either Mask R-CNN, DAFormer, or Mask2Former), we show the relative improvements in these models after implementing UDA techniques in Table 1. For the same synth-to-real setting, the UDA method proposed by Zhang improves the original model (w/o UDA) by +2.9 mAP and +5.9 mAP50. In contrast, our method achieves improvements of +6.7 mAP and +8.3 mAP50, respectively, demonstrating the superior effectiveness of the UDA approach in our framework. CVRN increases mAP50 from 16.2 to 25.9 (+9.7), and EDAPS increases it from 22.0 to 34.3 (+12.3). Our method, on the other hand, boosts it from 48.6 to 56.9 (+8.3). Even though our mAP50 gain is slightly lower, we work in a significantly higher performance regime, where improvement is harder to obtain.
4.2.2 Bidirectional Data Mixing.
To illustrate the advantages of the Bidirectional Data Mixing (BDM), we conduct experiments on different data mixing strategies without applying Rare-class Balancing (RcB) and Category Module Training (CMT). The results in the UrbanSyn Cityscapes setting are shown in Table 4.2.1. The mAP results reveal that both S2T and T2S mixing contribute to better performance. T2S is slightly inferior to S2T, which we believe is because the noise in the target pseudo-labels can inevitably cause interference in data mixing. Combining S2T and T2S mixing achieves the best result of 37.18 mAP, +5.23 higher than the plain model without mixing. Only applying S2T mixing neglects the instances in target images. Through BDM, we achieve a balance between learning target features and mitigating the negative impact of incorrect target pseudo-labels on the model. We also perform hyper-parameters searching for the weights of the BDM loss when applying all three UDA methods. The results are shown in Table 4.2.1. The best result yields at the setting of = 0.4 and = 0.6, however, the results are stable for the different assessed combinations.
@@
As mentioned, we perform CMT to introduce inductive bias, hoping to guide the model in better learning class features. To validate our conjecture, in the UrbanSyn Cityscapes setting, we perform CMT on two category groups, which are random compositions of classes (setting 2-5 in Table 5). Setting 1 is for full-class training. Overall, we observe that all four CMT models outperform the full-class training one. The best setting is 5, which achieves the highest improvement of +3.5 mAP compared to the model without CMT. We take this setting as the default for the rest of the experiments. However, three settings out of four show stable results. As mentioned in Sec. 3, we offer an alternative that enables the use of a single model during the testing phase. The category-free results for this approach are presented in Table 4.
| AP | AP50 | |||||||||||||||||||
| Model | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP | Diff. | Person | Rider | Car | Truck | Bus | Train | M.C | B.C | mAP50 | Diff. |
| UDA4Inst | 30.0 | 26.4 | 49.2 | 42.9 | 64.4 | 58.1 | 21.3 | 19.7 | 39.0 | - | 60.1 | 60.7 | 75.1 | 56.4 | 78.0 | 77.8 | 48.7 | 53.6 | 63.8 | - |
| category-free | 30.4 | 26.6 | 48.8 | 46.3 | 63.0 | 51.6 | 22.7 | 19.4 | 38.6 | -0.4 | 63.1 | 62.4 | 75.7 | 57.3 | 76.3 | 80.1 | 50.9 | 54.8 | 65.1 | +1.3 |
@@
To assess the impact of UDA4Inst components—BDM, RcB, and CMT—we perform an ablation study in the UrbanSyn Cityscapes setting in Table 6. The ablative models are named from M1 to M5 (from top to bottom) for illustration. Overall, M1 without UDA obtains the lowest mAP of 32.0, while M6, equipped with all three UDA methods, achieves the highest score of 39.0. When comparing M1 and M2, applying CMT can effectively improve the model performance by +2.8 mAP. The conclusion is further strengthened through the comparison of M3 and M4. The effect of BDM is demonstrated by the performance jump from 32.0 to 36.1 mAP (+4.1) from M1 to M3. Moreover, under the same circumstances as CMT, BDM helps to improve the precision in M4 across most classes compared to M2, contributing to a +2.9 mAP improvement. The RcB method, which is based on BDM and CMT, is applied in M5. Focusing on the defined rare classes, RcB helps M5 to obtain +10.8 and +1.8 AP gains to the M4 in ‘Train’ and ‘Motorcycle’, respectively. M4, using BDM alone without RcB, scores -1.3 mAP lower than M5. Comparing ablative models, BDM stands out as the most beneficial method for domain adaptation.
@@ 5 Conclusion
In this paper, we present a framework named UDA4Inst that aims to provide a strong baseline of synth-to-real UDA for instance segmentation. We incorporate three effective UDA methods into the vanilla Mask2Former training protocol: instance-level bidirectional data mixing, online rare-class balancing, and category module training. In a comprehensive evaluation of three synth-to-real domain settings: UrbanSyn/Synscapes/SYNTHIA Cityscapes, we show that UDA4Inst achieves significant improvements in instance segmentation compared to the source-only Mask2Former. This is the first work to provide competitive AP results on Synscapes and the recently published synthetic dataset UrbanSyn. Our proposed UDA4Inst model yields state-of-the-art results on SYNTHIA. We anticipate that this work will serve as a foundational framework for future research in UDA for instance segmentation.
5.0.1 Acknowledgements